# Audio Super-Resolution with Latent Bridge Models

### v2

@@ 1 Introduction
Audio super-resolution (SR) systems aim to generate the high-resolution (HR) audio waveform from the observed low-resolution (LR) waveform [47, 56, 109]. Audio SR plays a crucial roles in various applications, including historical recording restoration [73, 59, 74], hearing aids [30], and improvement of perceptual quality [62]. Previous works have explored audio SR with diverse methods such as mapping-based [60, 59], generative adversarial network (GAN)-based [69, 65], diffusion-based [72, 62], and recently proposed bridge-based [54, 45] methods. However, these efforts focus on the audio SR task in a limited scope, e.g., the speech signals in a benchmark dataset, without generalizing to a broader domain of different audio types. To extend audio SR to multiple domains spanning speech signals, sound effects, and music samples, Liu et al. [62] has recently explored diffusion models in the latent space of mel-spectrogram, enabling any-to-48 kHz audio SR. A2SB [45] has recently explored bridge models to synthesize the short-time Fourier transformation (STFT) representation, achieving any-to-44.1 kHz music SR.
Figure 1: AudioLBM significantly improves the perceptual quality of text-to-speech [111], text-to-audio [63], and text-to-music [53] generation, and outperforms the state-of-the-art any-to-48 kHz SR system AudioSR [62].
However, the quality of audio SR at scale is still limited. One major cause is the potential mismatch between their generative frameworks and the audio SR process. As shown in Fig. 2, AudioSR [62] synthesizes the missing part of the mel-spectrogram latent representation from an uninformative Gaussian prior, ignoring the fact that the LR waveform contains informative cues about the HR target. A2SB [45] formulates the audio SR analogous to image inpainting [58], where the removed regions are filled with Gaussian priors that remain uninformative about the target signal. Aiming for a high-quality audio SR system, we propose a latent bridge model (LBM) named AudioLBM, learning the LR-to-HR waveform SR process with a latent-to-latent generative framework. Specifically, we directly compress the audio waveform into a continuous latent representation, where the latent of the LR waveform demonstrates instructive information for the latent of the HR waveform and avoids the removed areas in previous works [41, 45]. Then, we design bridge models [14, 130] to enable a latent-to-latent generation process, which is inherently matched with the LR-to-HR waveform SR task and fully exploits the informative prior in the latent space.
@@
  • Zero-shot any-to-48 kHz audio SR evaluated on VCTK [118], Song-Describer-Dataset [68], and ESC-50 [79] demonstrate that our method outperforms prior systems with an average improvement of 21.5% in LSD [25] and 3.05% in ViSQOL [15] across audio and music, as well as achieves state-of-the-art objective and perceptual quality [87] for speech upsampling.
Figure 2: The top part shows how the low-resolution waveform is simulated during training via low-pass filtering, the middle part depicts the baseline method AudioSR [62] that synthesizes high-resolution content from Gaussian noise, and the bottom part presents overview of our proposed AudioLBM. It learns a latent-to-latent generation process between the low- and high-resolution waveform latent representations, namely $\bm{z_{x}}^{\text{LR}}\in\mathbb{R}^{c_{x}\times\frac{L}{r_{x}}}$ and $\bm{z_{x}}^{\text{HR}}\in\mathbb{R}^{c_{x}\times\frac{L}{r_{x}}}$, where $L$ is waveform length. $c_{x}$ and $r_{x}$ are the channel dimension and compression ratio of waveform latent. In contrast, AudioSR operates in latent space $\bm{z_{X}}\in\mathbb{R}^{c_{X}\times\frac{F}{r_{X}}\times\frac{T}{r_{X}}}$ of mel-spectrogram ${X}\in\mathbb{R}^{F\times T}$ with a noise-to-latent generation process, where $T$ and $F$ denote time and frequency bins of mel-spectrogram; $c_{X}$ and $r_{X}$ denote channel dimension and compression ratio of mel-spectrogram latent.
@@ 2 Related Work
Previous works on audio SR can be broadly categorized into spectral-based and waveform-based methods. Spectral-based methods, often referred to as bandwidth extension (BWE), formulate the task as a spectral inpainting problem in either time-frequency space [60, 59, 69, 96, 72, 41, 122, 126, 19, 46] or latent spectral space [62, 37, 110]. In contrast, waveform-based methods directly generate high-resolution signals in the time domain [47, 76, 49, 32, 125, 51, 54]. However, most existing works remain limited in specific domains such as speech or isolated music genres [73, 71], and are evaluated under fixed input resolutions [69, 65, 96], simplified degradation assumptions [32, 55], or limited output sampling rates [72, 46, 69]. These limitations hinder generalization to real-world scenarios involving diverse content and degradation types, such as polyphonic pop music that combines vocals, background instrumentation, and sound effects, or handling extremely low-resolution inputs. Moreover, performing SR beyond 48 kHz (e.g., 96 kHz and 192 kHz) remains unexplored, which provides further benefits in various professional applications, including mastering [57], post-production [7], spatial audio [8], and immersive content creation [101]. Although AudioSR [62] stands out as a scalable method achieving any-to-48 kHz SR across diverse domains, it suffers from two-stage compression pipeline and sub-optimal generative modeling paradigms for the SR task, leading to sub-optimal low-frequency fidelity, misalignment across frequency bands, and high-frequency artifacts [131], which ultimately limit perceptual SR quality. Therefore, building a unified and high-fidelity audio SR model that scales across speech signals, sound effects, and music samples remains a critical challenge. Detailed discussion of audio SR baseline methods is provided in Appendix H.2.2.
@@
Tractable bridge models [130, 14, 64, 52] have received increasing attention for enabling more effective and efficient data-to-data generation paradigms. Unlike diffusion models [34, 98] that generate data by reversing a Markov process that gradually injects noise into the input, bridge models begin from a deterministic and informative prior, and interpolate stochastically toward the target distribution. Recent studies have explored bridge models in various applications, including image translation [52, 103], image restoration [58, 18], dense prediction [38], image editing [11], quality assessment [128, 129] and text-to-speech synthesis [14]. Building on this paradigm, Bridge-SR [54] applies the Schrödinger Bridge (SB) framework directly in the waveform domain for speech SR, using a lightweight WaveNet-based score estimator [106, 44]. In parallel, A2SB [45] employs the SB formulation in the STFT domain for music bandwidth extension and inpainting using a U-Net backbone [89, 22]. However, directly modeling in the data space inherently limits generalization and poses challenges for stable training and flexible scaling to higher-resolution SR. These limitations motivate us to develop a more adaptive generation framework tailored to audio SR.
@@ 3.1 AudioLBM
Considering that the observed LR waveform has been an informative prior of the target HR waveform in the time-domain [54], we train a convolution-based variational autoencoder (VAE) [26] to directly compress the audio waveform into continuous representations, preserving the prior contained in the LR waveform for target generation. The details of VAE architecture and training configurations are introduced in Appendix B. Given the HR audio waveform $\bm{x}^{\text{HR}}\in\mathbb{R}^{L}$, we compute the LR counterpart $\bm{x}^{\text{LR}}\in\mathbb{R}^{L}$ with various low-pass filters to simulate the real-world degradations. Then, we compress them into latent representations $\bm{z}^{\text{HR}}\in\mathbb{R}^{c\times l}$ and $\bm{z}^{\text{LR}}\in\mathbb{R}^{c\times l}$ with a pre-trained encoder $\mathcal{E}(\bm{x})$, where $l=\frac{L}{r_{x}}$ and $r_{x}$ is the down-sampling factor, constructing the boundary distributions and establishing the bridge process as follows.
@@
Given $\bm{z}^{\text{LR}}\in\mathbb{R}^{c\times l}$ as the prior $\bm{z}_{T}$ at time step $t=T$ and $\bm{z}^{\text{HR}}\in\mathbb{R}^{c\times l}$ as the target $\bm{z}_{0}$ at time step $t=0$, we build a bridge process [14] to connect the boundary distributions (see details in Appendix C):
@@
There exist multiple equivalent training objectives for bridge models. We empirically find that in the latent space, our AudioLBM achieves higher synthesis quality by using a noise predictor, compared to the data predictor used by recent bridge-related works in waveform [54] or mel-spectrogram domain [14]. In this regard, we formulate the denoising objective as:
@@
The forward process of bridge models has a reverse process sharing the same marginal distribution $p_{t}(\bm{z}_{t}|\bm{z}_{0},\bm{z}_{T})$ [14, 130]. For signal generation, starting from the prior $\bm{z}^{\text{LR}}$, we use a first-order SDE-based sampler. From the time step $s$ to the time step $t\in\left[0,s\right)$, the first-order discretization gives:
@@
As shown in Fig. 2, the diffusion-based audio upsampling employs a conditional noise-to-latent sampling trajectory, where the prior provides limited information for the target. In comparison, our method employs a latent-to-latent sampling trajectory from $\bm{z}^{\text{LR}}$ to $\bm{z}^{\text{HR}}$, which has been aligned with the LR-to-HR audio upsampling. Moreover, in the continuous latent space directly compressed from the audio waveform, the representation of LR waveform provides instructive information for the generation target, rather than suffering from area removal in a latent space compressed by the STFT representation [45] or mel-spectrogram [122].
@@ 3.2 Frequency-aware LBMs
Compared to most text-to-speech [111, 24] or text-to-audio [61, 63, 27, 50] generation systems that synthesize the audio samples at a sampling rate below $24$ or $44.1$ kHz, it is more expensive to collect the audio samples at a sampling rate of $48$ kHz for training audio up-sampling models. To address the limitation of dataset scale, we propose frequency-aware LBMs, enabling an any-to-any upsampling process at the training stage. Specifically, given an audio signal $\bm{x}\in\mathbb{R}^{L}$, we first filter an HR waveform $\bm{x}^{\text{HR}}\in\mathbb{R}^{L}$ with a sampling rate $\text{SR}_{\bm{x}^{\text{HR}}}$ [^1] To avoid ambiguity, we define the low-pass filter (LPF) by its sampling rate, corresponding to filtering with cutoff at its Nyquist frequency. that is lower than the sampling rate of $\bm{x}$, while it has already provided the frequency band where the audio information predominantly concentrates (detailed in Appendix C). Then, we compute the LR version $\bm{x}^{\text{LR}}\in\mathbb{R}^{L}$ with a sampling rate $\text{SR}_{\bm{x}^{\text{LR}}}$ uniformly sampled from $\mathcal{U}\left(0,\text{SR}_{\bm{x}^{\text{HR}}}\right)$, constructing a new LR-HR data pair rather than using a fixed sampling rate for $\bm{x}^{\text{HR}}$.
At each training iteration, we first compress the $\bm{x}^{\text{HR}}$ and $\bm{x}^{\text{LR}}$ into latent $\bm{z}^{\text{HR}}$ and $\bm{z}^{\text{LR}}$, and then take their sampling rate $\text{SR}_{\bm{x}^{\text{HR}}}$ and $\text{SR}_{\bm{x}^{\text{LR}}}$ as model input, encouraging LBMs to explicitly learn an any-to-any upsampling process. Specifically, we extract a sinusoidal embedding of quantized $f_{\text{target}}=\text{Quantize}(\text{SR}_{\bm{x}^{\text{HR}}}/2)$ and continuous $f_{\text{prior}}=\text{SR}_{\bm{x}^{\text{LR}}}/2$ and prepend them as two additional tokens of our DiT [78, 4]-based noise prediction network, expanding the feature space to $\mathbb{R}^{c\times(l+2)}$. Towards stronger training performance, we leverage the constant scaling factor $s$ for training bridge models in the waveform domain [54], rescaling the latent with $\tilde{\bm{z}}=s*\bm{z}$ for stable training, leading to the training objective of our frequency-aware LBMs:
@@ 3.3 Audio SR beyond 48 kHz
Figure 3: LBMs can be naturally extended into higher-resolution waveform generation with a cascaded paradigm, where prior argumentation is utilized to avoid cascading artifacts and accumulating errors between stages.
@@
Beyond 48 kHz, the audio at an ultra-high sampling rate (e.g., 96 kHz or 192 kHz) provides engineering advantages and post-processing flexibility [85]. However, scaling the upsampling system to such a sampling rate is constrained by the limited capacity of a single model and the scarcity of high-resolution training data, which has not been addressed to date. Here, we propose an extended version of AudioLBM that progressively improves the reconstructed signals from 48 kHz to higher sampling rates. Taking the upsampling to 96 kHz as an example, we introduce cascaded LBMs. We first train a VAE-based compression network $\mathcal{E}(\bm{x}^{\text{UHR}})$ for the ultra HR audio waveform $\bm{x}^{\text{UHR}}$, enabling us to design a new AudioLBM $\bm{\epsilon}^{\text{UHR}}_{\theta}$ in the latent space $\bm{z}^{\text{UHR}}$. At generation stage, the waveform $\bm{x}^{\text{LR}}$ is first upsampled to $\bm{x}^{\text{HR}}$ with the first-stage LBM, and then taken as the input of the second-stage model, further upsampled to $\bm{x}^{\text{UHR}}$ with a latent-to-latent process from $\bm{z}^{\text{HR}}$ to $\bm{z}^{\text{UHR}}$ with Eq. (3), as shown in Figure 3. In the scenario of 192 kHz upsampling, a third-stage LBM is cascaded. However, as the data has been extremely expensive to collect, which restricts the model performance, we propose a fine-tuning method to leverage the pretrained LBM at the second stage, improving the synthesis quality.
@@
One of the potential limitations of cascaded generative models is the cascading errors in inference [35]. In cascaded LBMs, the LBM at the first stage approximates the ground-truth distribution of $\bm{x}^{\text{HR}}$, while we inevitably suffer from a mismatch between the synthesized data and the GT signal because of the imperfect learning of $\bm{\epsilon}_{\theta}$. Hence, we propose two prior augmentation strategies in the waveform and latent space, respectively, for the cascaded bridge models $\bm{\epsilon}^{\text{UHR}}_{\theta}$, alleviating the mismatch issue with the first-stage LBM $\bm{\epsilon}_{\theta}$ and therefore strengthening the upsampling performance. Firstly, considering the difficulty for generative models to reconstruct fine-grained waveform features, we introduce a degradation operation in the waveform domain. Specifically, in the second upsampling stage, we randomly remove a small portion of high-frequency details near the Nyquist boundary of the prior and obtain a low-pass-filtered waveform, denoted as $\text{LPF}(\bm{x}^{\text{HR}})$. Furthermore, solely simulating the previous stage’s output via waveform degradation is insufficient to address potential large-scale artifacts, which may be reflected in unclear harmonics or imbalanced energy distribution. To mitigate this, we further introduce latent-space blurring, which perturbs the audio by applying dynamic Gaussian smoothing to its latent representation along the time axis. Unlike previous diffusion-based [35] or flow-based [95] methods where the latent is corrupted with random noise for argumentation, the blurring strategy provides a deterministic degradation that aligns with the Dirac boundary distributions in bridge models [14], leading to the following training objective for the cascaded bridge model $\bm{\epsilon}^{\text{UHR}}_{\theta}$:

### v3

@@ 1 Introduction
Audio super-resolution (SR) systems aim to generate the high-resolution (HR) audio waveform from the observed low-resolution (LR) waveform [kuleshov2017audio, lim2018time, wang2021towards]. Audio SR plays a crucial roles in various applications, including historical recording restoration [moliner2024blind, liu2021voicefixer, moliner2024diffusion], hearing aids [fullgrabe2010preliminary], and improvement of perceptual quality [liu2024audiosr]. Previous works have explored audio SR with diverse methods such as mapping-based [liu2022neural, liu2021voicefixer], generative adversarial network (GAN)-based [mandel2023aero, lu2024towards], diffusion-based [moliner2023solving, liu2024audiosr], and recently proposed bridge-based [li2025bridge, kong2025a2sb] methods. However, these efforts focus on the audio SR task in a limited scope, e.g., the speech signals in a benchmark dataset, without generalizing to a broader domain of different audio types. To extend audio SR to multiple domains spanning speech signals, sound effects, and music samples, liu2024audiosr has recently explored diffusion models in the latent space of mel-spectrogram, enabling any-to-48 kHz audio SR. A2SB [kong2025a2sb] has recently explored bridge models to synthesize the short-time Fourier transformation (STFT) representation, achieving any-to-44.1 kHz music SR.
Figure 1: AudioLBM significantly improves the perceptual quality of text-to-speech [wang2024maskgct], text-to-audio [liu2024audioldm], and text-to-music [li2024quality] generation, and outperforms the state-of-the-art any-to-48 kHz SR system AudioSR [liu2024audiosr].
However, the quality of audio SR at scale is still limited. One major cause is the potential mismatch between their generative frameworks and the audio SR process. As shown in Fig. 2, AudioSR [liu2024audiosr] synthesizes the missing part of the mel-spectrogram latent representation from an uninformative Gaussian prior, ignoring the fact that the LR waveform contains informative cues about the HR target. A2SB [kong2025a2sb] formulates the audio SR analogous to image inpainting [liu2023i2sb], where the removed regions are filled with Gaussian priors that remain uninformative about the target signal. Aiming for a high-quality audio SR system, we propose a latent bridge model (LBM) named AudioLBM, learning the LR-to-HR waveform SR process with a latent-to-latent generative framework. Specifically, we directly compress the audio waveform into a continuous latent representation, where the latent of the LR waveform demonstrates instructive information for the latent of the HR waveform and avoids the removed areas in previous works [kim2024audio, kong2025a2sb]. Then, we design bridge models [chen2023schrodinger, zhou2023denoising] to enable a latent-to-latent generation process, which is inherently matched with the LR-to-HR waveform SR task and fully exploits the informative prior in the latent space.
@@
  • Zero-shot any-to-48 kHz audio SR evaluated on VCTK [yamagishi2019cstr], Song-Describer-Dataset [manco2023song], and ESC-50 [piczak2015esc] demonstrate that our method outperforms prior systems with an average improvement of 21.5% in LSD [erell1990estimation] and 3.05% in ViSQOL [chinen2020visqol] across audio and music, as well as achieves state-of-the-art objective and perceptual quality [ristea2025icassp] for speech upsampling.
Figure 2: The top part shows how the low-resolution waveform is simulated during training via low-pass filtering, the middle part depicts the baseline method AudioSR [liu2024audiosr] that synthesizes high-resolution content from Gaussian noise, and the bottom part presents overview of our proposed AudioLBM. It learns a latent-to-latent generation process between the low- and high-resolution waveform latent representations, namely $\bm{z_{x}}^{\text{LR}}\in\mathbb{R}^{c_{x}\times\frac{L}{r_{x}}}$ and $\bm{z_{x}}^{\text{HR}}\in\mathbb{R}^{c_{x}\times\frac{L}{r_{x}}}$, where $L$ is waveform length. $c_{x}$ and $r_{x}$ are the channel dimension and compression ratio of waveform latent. In contrast, AudioSR operates in latent space $\bm{z_{X}}\in\mathbb{R}^{c_{X}\times\frac{F}{r_{X}}\times\frac{T}{r_{X}}}$ of mel-spectrogram ${X}\in\mathbb{R}^{F\times T}$ with a noise-to-latent generation process, where $T$ and $F$ denote time and frequency bins of mel-spectrogram; $c_{X}$ and $r_{X}$ denote channel dimension and compression ratio of mel-spectrogram latent.
@@ 2 Related Work
Previous works on audio SR can be broadly categorized into spectral-based and waveform-based methods. Spectral-based methods, often referred to as bandwidth extension (BWE), formulate the task as a spectral inpainting problem in either time-frequency space [liu2022neural, liu2021voicefixer, mandel2023aero, shuai2023mdctgan, moliner2023solving, kim2024audio, yun2025flowhigh, zhang2025vm, comunitaspecmaskgit, ku2025generative] or latent spectral space [liu2024audiosr, im2025flashsr, wang2023audit]. In contrast, waveform-based methods directly generate high-resolution signals in the time domain [kuleshov2017audio, nguyen2022tunet, lee2021nu, han2022nu, zhang2021wsrglow, lee2025wave, li2025bridge]. However, most existing works remain limited in specific domains such as speech or isolated music genres [moliner2024blind, moliner2022behm], and are evaluated under fixed input resolutions [mandel2023aero, lu2024towards, shuai2023mdctgan], simplified degradation assumptions [han2022nu, li2025apollo], or limited output sampling rates [moliner2023solving, ku2025generative, mandel2023aero]. These limitations hinder generalization to real-world scenarios involving diverse content and degradation types, such as polyphonic pop music that combines vocals, background instrumentation, and sound effects, or handling extremely low-resolution inputs. Moreover, performing SR beyond 48 kHz (e.g., 96 kHz and 192 kHz) remains unexplored, which provides further benefits in various professional applications, including mastering [link1999digital], post-production [black1999anti], spatial audio [blauert1997spatial], and immersive content creation [stuart1997coding]. Although AudioSR [liu2024audiosr] stands out as a scalable method achieving any-to-48 kHz SR across diverse domains, it suffers from two-stage compression pipeline and sub-optimal generative modeling paradigms for the SR task, leading to sub-optimal low-frequency fidelity, misalignment across frequency bands, and high-frequency artifacts [zhu2024musichifi], which ultimately limit perceptual SR quality. Therefore, building a unified and high-fidelity audio SR model that scales across speech signals, sound effects, and music samples remains a critical challenge. Detailed discussion of audio SR baseline methods is provided in Appendix H.2.2.
@@
Tractable bridge models [zhou2023denoising, chen2023schrodinger, liu2023learning, li2023bbdm] have received increasing attention for enabling more effective and efficient data-to-data generation paradigms. Unlike diffusion models [ho2020denoising, song2020score] that generate data by reversing a Markov process that gradually injects noise into the input, bridge models begin from a deterministic and informative prior, and interpolate stochastically toward the target distribution. Recent studies have explored bridge models in various applications, including image translation [li2023bbdm, su2022dual], image restoration [liu2023i2sb, chung2023direct], dense prediction [ji2024dpbridge], image editing [chadebec2025lbm], quality assessment [zhou2025asal, zhou25phi] and text-to-speech synthesis [chen2023schrodinger]. Building on this paradigm, Bridge-SR [li2025bridge] applies the Schrödinger Bridge (SB) framework directly in the waveform domain for speech SR, using a lightweight WaveNet-based score estimator [van2016wavenet, kong2020diffwave]. In parallel, A2SB [kong2025a2sb] employs the SB formulation in the STFT domain for music bandwidth extension and inpainting using a U-Net backbone [ronneberger2015u, dhariwal2021diffusion]. However, directly modeling in the data space inherently limits generalization and poses challenges for stable training and flexible scaling to higher-resolution SR. These limitations motivate us to develop a more adaptive generation framework tailored to audio SR.
@@ 3.1 AudioLBM
Considering that the observed LR waveform has been an informative prior of the target HR waveform in the time-domain [li2025bridge], we train a convolution-based variational autoencoder (VAE) [evanslong] to directly compress the audio waveform into continuous representations, preserving the prior contained in the LR waveform for target generation. The details of VAE architecture and training configurations are introduced in Appendix B. Given the HR audio waveform $\bm{x}^{\text{HR}}\in\mathbb{R}^{L}$, we compute the LR counterpart $\bm{x}^{\text{LR}}\in\mathbb{R}^{L}$ with various low-pass filters to simulate the real-world degradations. Then, we compress them into latent representations $\bm{z}^{\text{HR}}\in\mathbb{R}^{c\times l}$ and $\bm{z}^{\text{LR}}\in\mathbb{R}^{c\times l}$ with a pre-trained encoder $\mathcal{E}(\bm{x})$, where $l=\frac{L}{r_{x}}$ and $r_{x}$ is the down-sampling factor, constructing the boundary distributions and establishing the bridge process as follows.
@@
Given $\bm{z}^{\text{LR}}\in\mathbb{R}^{c\times l}$ as the prior $\bm{z}_{T}$ at time step $t=T$ and $\bm{z}^{\text{HR}}\in\mathbb{R}^{c\times l}$ as the target $\bm{z}_{0}$ at time step $t=0$, we build a bridge process [chen2023schrodinger] to connect the boundary distributions (see details in Appendix C):
@@
There exist multiple equivalent training objectives for bridge models. We empirically find that in the latent space, our AudioLBM achieves higher synthesis quality by using a noise predictor, compared to the data predictor used by recent bridge-related works in waveform [li2025bridge] or mel-spectrogram domain [chen2023schrodinger]. In this regard, we formulate the denoising objective as:
@@
The forward process of bridge models has a reverse process sharing the same marginal distribution $p_{t}(\bm{z}_{t}|\bm{z}_{0},\bm{z}_{T})$ [chen2023schrodinger, zhou2023denoising]. For signal generation, starting from the prior $\bm{z}^{\text{LR}}$, we use a first-order SDE-based sampler. From the time step $s$ to the time step $t\in\left[0,s\right)$, the first-order discretization gives:
@@
As shown in Fig. 2, the diffusion-based audio upsampling employs a conditional noise-to-latent sampling trajectory, where the prior provides limited information for the target. In comparison, our method employs a latent-to-latent sampling trajectory from $\bm{z}^{\text{LR}}$ to $\bm{z}^{\text{HR}}$, which has been aligned with the LR-to-HR audio upsampling. Moreover, in the continuous latent space directly compressed from the audio waveform, the representation of LR waveform provides instructive information for the generation target, rather than suffering from area removal in a latent space compressed by the STFT representation [kong2025a2sb] or mel-spectrogram [yun2025flowhigh].
@@ 3.2 Frequency-aware LBMs
Compared to most text-to-speech [wang2024maskgct, du2024cosyvoice] or text-to-audio [liu2023audioldm, liu2024audioldm, evans2024fast, lee2024etta] generation systems that synthesize the audio samples at a sampling rate below $24$ or $44.1$ kHz, it is more expensive to collect the audio samples at a sampling rate of $48$ kHz for training audio up-sampling models. To address the limitation of dataset scale, we propose frequency-aware LBMs, enabling an any-to-any upsampling process at the training stage. Specifically, given an audio signal $\bm{x}\in\mathbb{R}^{L}$, we first filter an HR waveform $\bm{x}^{\text{HR}}\in\mathbb{R}^{L}$ with a sampling rate $\text{SR}_{\bm{x}^{\text{HR}}}$ [^1] To avoid ambiguity, we define the low-pass filter (LPF) by its sampling rate, corresponding to filtering with cutoff at its Nyquist frequency. that is lower than the sampling rate of $\bm{x}$, while it has already provided the frequency band where the audio information predominantly concentrates (detailed in Appendix C). Then, we compute the LR version $\bm{x}^{\text{LR}}\in\mathbb{R}^{L}$ with a sampling rate $\text{SR}_{\bm{x}^{\text{LR}}}$ uniformly sampled from $\mathcal{U}\left(0,\text{SR}_{\bm{x}^{\text{HR}}}\right)$, constructing a new LR-HR data pair rather than using a fixed sampling rate for $\bm{x}^{\text{HR}}$.
At each training iteration, we first compress the $\bm{x}^{\text{HR}}$ and $\bm{x}^{\text{LR}}$ into latent $\bm{z}^{\text{HR}}$ and $\bm{z}^{\text{LR}}$, and then take their sampling rate $\text{SR}_{\bm{x}^{\text{HR}}}$ and $\text{SR}_{\bm{x}^{\text{LR}}}$ as model input, encouraging LBMs to explicitly learn an any-to-any upsampling process. Specifically, we extract a sinusoidal embedding of quantized $f_{\text{target}}=\text{Quantize}(\text{SR}_{\bm{x}^{\text{HR}}}/2)$ and continuous $f_{\text{prior}}=\text{SR}_{\bm{x}^{\text{LR}}}/2$ and prepend them as two additional tokens of our DiT [peebles2023scalable, bao2023all]-based noise prediction network, expanding the feature space to $\mathbb{R}^{c\times(l+2)}$. Towards stronger training performance, we leverage the constant scaling factor $s$ for training bridge models in the waveform domain [li2025bridge], rescaling the latent with $\tilde{\bm{z}}=s*\bm{z}$ for stable training, leading to the training objective of our frequency-aware LBMs:
@@ 3.3 Audio SR beyond 48 kHz
Figure 3: LBMs can be naturally extended into higher-resolution waveform generation with a cascaded paradigm, where prior augmentation is utilized to avoid cascading artifacts and accumulating errors between stages.
@@
Beyond 48 kHz, the audio at an ultra-high sampling rate (e.g., 96 kHz or 192 kHz) provides engineering advantages and post-processing flexibility [reiss2016meta]. However, scaling the upsampling system to such a sampling rate is constrained by the limited capacity of a single model and the scarcity of high-resolution training data, which has not been addressed to date. Here, we propose an extended version of AudioLBM that progressively improves the reconstructed signals from 48 kHz to higher sampling rates. Taking the upsampling to 96 kHz as an example, we introduce cascaded LBMs. We first train a VAE-based compression network $\mathcal{E}(\bm{x}^{\text{UHR}})$ for the ultra HR audio waveform $\bm{x}^{\text{UHR}}$, enabling us to design a new AudioLBM $\bm{\epsilon}^{\text{UHR}}_{\theta}$ in the latent space $\bm{z}^{\text{UHR}}$. At generation stage, the waveform $\bm{x}^{\text{LR}}$ is first upsampled to $\bm{x}^{\text{HR}}$ with the first-stage LBM, and then taken as the input of the second-stage model, further upsampled to $\bm{x}^{\text{UHR}}$ with a latent-to-latent process from $\bm{z}^{\text{HR}}$ to $\bm{z}^{\text{UHR}}$ with Eq. (3), as shown in Figure 3. In the scenario of 192 kHz upsampling, a third-stage LBM is cascaded. However, as the data has been extremely expensive to collect, which restricts the model performance, we propose a fine-tuning method to leverage the pretrained LBM at the second stage, improving the synthesis quality.
@@
One of the potential limitations of cascaded generative models is the cascading errors in inference [ho2022cascaded]. In cascaded LBMs, the LBM at the first stage approximates the ground-truth distribution of $\bm{x}^{\text{HR}}$, while we inevitably suffer from a mismatch between the synthesized data and the GT signal because of the imperfect learning of $\bm{\epsilon}_{\theta}$. Hence, we propose two prior augmentation strategies in the waveform and latent space, respectively, for the cascaded bridge models $\bm{\epsilon}^{\text{UHR}}_{\theta}$, alleviating the mismatch issue with the first-stage LBM $\bm{\epsilon}_{\theta}$ and therefore strengthening the upsampling performance. Firstly, considering the difficulty for generative models to reconstruct fine-grained waveform features, we introduce a degradation operation in the waveform domain. Specifically, in the second upsampling stage, we randomly remove a small portion of high-frequency details near the Nyquist boundary of the prior and obtain a low-pass-filtered waveform, denoted as $\text{LPF}(\bm{x}^{\text{HR}})$. Furthermore, solely simulating the previous stage’s output via waveform degradation is insufficient to address potential large-scale artifacts, which may be reflected in unclear harmonics or imbalanced energy distribution. To mitigate this, we further introduce latent-space blurring, which perturbs the audio by applying dynamic Gaussian smoothing to its latent representation along the time axis. Unlike previous diffusion-based [ho2022cascaded] or flow-based [schusterbauer2024fmboost] methods where the latent is corrupted with random noise for augmentation, the blurring strategy provides a deterministic degradation that aligns with the Dirac boundary distributions in bridge models [chen2023schrodinger], leading to the following training objective for the cascaded bridge model $\bm{\epsilon}^{\text{UHR}}_{\theta}$: